OpenStudy (acespeedfighter):

The solution to 7x = 42 is x = ___. (3 points) 4 5 6 7

OpenStudy (acespeedfighter):

I think its 6

OpenStudy (legomyego180):

whats 7 times 6

alones (alones):

Divide both sides by 7. \[\frac{ 7x }{ 7 }=\frac{ 42 }{ ? }\]

alones (alones):

lolol i mean \(\frac{ 42 }{ 7}\) and i think you're correct ;)